### v1

@@ 4 The general case
In this section we provide a generalization to the geodesic setting of lower bounds that are known for the original problem in the plane. In particular, we generalize results from [13, 14, 18]. We note that upper bounds for the Euclidean setting are in general also valid for the geodesic setting, since one can take the constructions provided in [2, 11, 13], and enclose the relevant set of points in a bounding polygon, big enough so that the geodesic disks through every three points coincide with the corresponding Euclidean disks.
We begin by adapting some definitions and results from [18] to our setting. The intersection number $I(S)$ of a set of points $S$ in a polygon is defined as the number of different geodesic paths $g(u,v)$, $g(x,y)$ for $u,v,x,y\in S$ such that $g(u,v)\cap g(x,y)\neq\emptyset$. Then, the intersection graph $G(S)=(V(G(S)),E(G(S)))$ is defined with $V(G(S))=\{g(u,v)|u,v\in S,u\neq v\}$, and two vertices $g(u,v)$, $g(x,y)$ being joined by an edge in $G(S)$ if $g(u,v)\cap g(x,y)\neq\emptyset$. Finally, the oriented graph $\overrightarrow{G}(S)$ of $G(S)$ is defined orienting an edge $g(u,v)-g(x,y)$ as $g(u,v)\rightarrow g(x,y)$ if any geodesic disk through $u$ and $v$ contains $x$ or $y$, or otherwise orienting $g(x,y)\rightarrow g(u,v)$. This orientation is consistent because of Lemma 2.
@@
For any set $S$ of $n\geq 5$ points in a polygon, $\Pi(n)\geq\left\lceil\frac{n-2}{60}\right\rceil\approx 0.0166n$.
@@
By Kuratowski’s Theorem [16], each subset $S^{\prime}\subset S$, with $|S^{\prime}|=5$, contains four points $w,x,y,z$ such that $g(w,x)\cap g(y,z)\neq\emptyset$. Moreover, the subset $\{w,x,y,z\}$ appears in exactly $n-4$ subsets of $S$ with five elements. Thus, $I(S)\geq\frac{\binom{n}{5}}{n-4}$. By the definition of the intersection graph $G(S)$, $|E(G(S))|=I(S)$, thus $|E(G(S))|\geq\frac{\binom{n}{5}}{n-4}$.
@@
$\sum_{g(x,y)\in V(\overrightarrow{G}(S))}d^{+}(g(x,y))=|E(G(S))|\geq\frac{\binom{n}{5}}{n-4}.$
@@
This means that any geodesic disk through $u_{0}$ and $v_{0}$ contains at least one endpoint of $\frac{(n-2)(n-3)}{60}$ geodesic paths $g(x,y),\ x,y\in S$. Since each point could appear in at most $n-3$ of these geodesic paths, any geodesic disk through $u_{0}$ and $v_{0}$ contains at least $\left\lceil\frac{\frac{(n-2)(n-3)}{60}}{n-3}\right\rceil=\left\lceil\frac{n-2}{60}\right\rceil$ points of $S$ different from $u_{0}$ and $v_{0}$. ∎
@@
Among any set $S^{\prime}$ of $k$ points in a polygon $P$, there are at least $\binom{k-3}{2}$ pairs of points $x,y\in S^{\prime}$ such that every geodesic disk through $x$ and $y$ contains at least one of the other $k-2$ points of $S^{\prime}$.
@@
Let $G=(V,E)$ be the graph where $V\!=\!S^{\prime}$, and two vertices are joined by an edge if at least one of the geodesic disks through them does not contain any of the other $k-2$ points of $S^{\prime}$. The edge between $u$ and $v$ is drawn as the geodesic path $g(u,v)$. We claim that $G$ is planar, and the edges can be drawn without crossings inside $P$. Suppose there are two edges $g(u,v),g(x,y)$ in $G$, for $u,v,x,y\in S^{\prime}$, that intersect. By Lemma 2, one of the edges has the property that every geodesic disk through both of its endpoints contains at least one of the endpoints of the other diagonal, which contradicts the definition of the edges of $G$. Thus, $G$ is a planar graph with $k$ vertices. By Euler’s formula, $|E|\leq 3k-6$, so $S^{\prime}$ contains at least $\binom{k}{2}-(3k-6)=\binom{k-3}{2}$ pairs that satisfy the property. ∎
@@
For any set $S$ of $n\geq 8$ points in a polygon, $\Pi(n)\geq\left\lceil\frac{5}{84}(n-2)\right\rceil=\left\lceil\frac{n-2}{16.8}\right\rceil\approx 0.0595n$.
@@
Two points in a subset $S^{\prime}\subset S$ of $k$ points dominate $S^{\prime}$ if every geodesic disk through them contains at least one of the other $k-2$ points of $S^{\prime}$. There are $\binom{n}{k}$ subsets of size $k$. By Lemma 6, each of these sets has $\binom{k-3}{2}$ dominating pairs. The average number of subsets dominated by a pair is $\frac{\binom{k-3}{2}\binom{n}{k}}{\binom{n}{2}}$. Thus, there is a pair of points $\{u,v\}\in S^{\prime}$ that dominates at least $\frac{\binom{k-3}{2}\binom{n}{k}}{\binom{n}{2}}$ subsets of $S$.
For each of these subsets, every geodesic disk through $u$ and $v$ contains at least one of the other $k-2$ points of $S^{\prime}$. Since a point of $S\setminus\{u,v\}$ can be in at most $\binom{n-3}{k-3}$ subsets that include $u$ and $v$, every geodesic disk through $u$ and $v$ contains at least
@@
other points of $S$. The largest coefficient of $n-2$ is $\frac{5}{84}$, obtained for $k=8$ and for $k=9$. ∎
We now present our strongest lower bound for $\Pi(n)$ by generalizing the approach by Edelsbrunner et al. [12] to the geodesic setting.
@@
For any set $S$ of $n$ points in a polygon, $\Pi(n)\geq\left\lceil\frac{n}{5}\right\rceil+1$.
@@
Unless specified otherwise, in this proof, order-$k$ diagrams refer to order-$k$ geodesic Voronoi diagrams. Recall that a geodesic disk through two points containing $k-1$ points in its interior has its center on an edge of the order-$k$ diagram. Consider the set of all disks that pass through two points $p,q\in S$ in the polygon. These disks are all centered on the bisector $b(p,q)$, and there are at most $n-2$ points on the bisector where a third point of $S$ is on the boundary of the disk. This partitions the bisector into at most $n-1$ pieces such that any circle through $p$ and $q$ centered on any point belonging to a fixed piece $s$ contains in its interior the same number of points $\rho(s)$. The pieces $s$ such that $\rho(s)=k-1$ are precisely the edges of the order-$k$ diagram.
@@
Now consider some fixed bisector $b(p,q)$. The edges of the order-$1$ through order-$k$ diagrams on $b(p,q)$ form a number of connected components (which is strictly less than the total number of edges of all the diagrams). Let $\lambda_{k}$ be the number of connected components formed by all the bisectors of the order-$1$ through order-$k$ diagrams. The goal is to find the largest value of $k$ such that $\lambda_{k}<\binom{n}{2}$. Every bisector is completely covered by a single connected component when $k=n-2$. When $\lambda_{k}<\binom{n}{2}$, by the pigeonhole principle there exists a bisector $b(p,q)$ that does not have any edges of the order-$1$ through order-$k$ diagrams. Therefore, every geodesic disk with $p$ and $q$ on its boundary must contain in its interior at least $k$ points.
@@
We derive a formula for $\lambda_{k}$ in terms of the number of edges in the order-$1$ to the order-$k$ diagram by considering how the pieces must “pair up”, in the sense that a piece $s$ with $\rho(s)=i$ can only be adjacent to pieces $s^{\prime}$ with $\rho(s^{\prime})=i\pm 1$. If a piece $s$ with $\rho(s)=k$ is adjacent to two pieces $s_{1}$ and $s_{2}$ with $\rho(s_{1})=\rho(s_{2})=k-1$, the number of connected components is decreased by one. If $s$ is adjacent to only one piece $s_{1}$ with $\rho(s_{1})=k-1$, it extends an existing component. Otherwise, if it is not adjacent to any piece $s_{1}$ with $\rho(s_{1})=k-1$, it creates a new connected component. Let $\epsilon_{a},\epsilon_{b}$ and $\epsilon_{c}$ be, respectively, the number of pieces $s$ with $\rho(s)=k$ of the first, second, and third type. Then we have
$\lambda_{k}=\lambda_{k-1}+\epsilon_{c}-\epsilon_{a}$ (1)
@@
Let $e_{k}$ be the number of edges in an order-$k$ diagram and let $S_{k}$ be the number of edges of the order-$k$ diagram that intersect the boundary. The total number of endpoints of all order-$k$ edges is given by $2e_{k}-S_{k}$ since the edges that intersect the boundary of the polygon only contribute one endpoint. The edges intersecting the boundary are analogous to the unbounded edges in the Euclidean setting. Moreover, we note that $e_{k}=\epsilon_{a}+\epsilon_{b}+\epsilon_{c}$.
@@
Let $\phi_{i}$ be the number of unpaired endpoints in $\lambda_{i}$. Consider all the unpaired endpoints of the connected components of $\lambda_{k-1}$. These endpoints can only come from edges in the order-${k-1}$ diagram since all edges of the order-$i$ diagrams for $i\in\{1,\ldots,k-2\}$ are all paired. Therefore, these unpaired endpoints can only pair up with edges from the order-$k$ diagram, which implies that $\phi_{k-1}=2\epsilon_{a}+\epsilon_{b}$. Thus, from Equation (1) and the fact that $e_{k}=\epsilon_{a}+\epsilon_{b}+\epsilon_{c}$, we have that $\lambda_{k}=\lambda_{k-1}+e_{k}-\phi_{k-1}$. If we iterate over $k$, we get that $\lambda_{k}=(e_{k}-\phi_{k-1})+(e_{k-1}-\phi_{k-2})+\ldots+(e_{2}-\phi_{1})+e_{1}=\sum_{i=1}^{k}e_{i}-\sum_{i=1}^{k-1}\phi_{i}$. Note that $\phi_{i-1}$ is the number of unpaired endpoints in $\lambda_{i-1}$, which are the paired endpoints of $\lambda_{i}$. Therefore, $\phi_{i}+\phi_{i-1}=2e_{i}-S_{i}$–the number of endpoints in order-$i$ diagrams. We use this to rewrite the equation for $\lambda_{k}$ in terms of $e_{i}$ and $S_{i}$.
$\lambda_{k}$ $=\sum_{i=1}^{k}e_{i}-[(\phi_{k-1}+\phi_{k-2})+(\phi_{k-3}+\phi_{k-4})+\ldots]\mbox{ (end of sum depends on parity of $k$)}$
@@
$=\sum_{i=1}^{k}(-1)^{k-i}e_{i}+\left(S_{k-1}+S_{k-3}+\ldots+S_{2\mbox{-}or\mbox{-}1}\right)\mbox{ ($S_{2}$ or $S_{1}$ depending on parity of $k$)}$
@@
We use the fact that $\sum_{i=1}^{j}S_{i}\geq j(j+1)$, see [8, Lemma $11$], which implies that
@@
which is a quadratic polynomial in $k$. We note that $\lambda_{k}<\binom{n}{2}$ whenever $k<\frac{n}{5}$, i.e., $k\leq\left\lceil\frac{n}{5}\right\rceil-1$. Then, as mentioned above, the pigeonhole principle implies that every geodesic disk with $p$ and $q$ on its boundary must contain in its interior at least $k\leq\left\lceil\frac{n}{5}\right\rceil-1$ points. Thus, we get that $\Pi(n)\geq\left\lceil\frac{n}{5}\right\rceil+1$, with the plus one accounting for $p$ and $q$. ∎
@@ 4.1 A Constant Fraction of Points Inside and Outside Geodesic Disks
In this section we show that for a point set $S$, there is a pair of points $x,y\in S$ such that any disk through them contains a constant fraction of points of $S$, both inside and outside the disks. We obtain this bound by modifying the previous proof of Theorem 2 by using Lemma 7. For completeness, we add here the proof of the bound.
@@
Among any set $S^{\prime}$ of $k$ points in a polygon, there are at least $\binom{k-3}{2}$ pairs of points $x,y\in S^{\prime}$ such that every geodesic disk through $x$ and $y$ leaves outside at least one of the other $k-2$ points of $S^{\prime}$.
@@
Let $G=(V,E)$ be the graph where $V\!=\!S^{\prime}$, and two vertices are joined by an edge if at least one of the geodesic disks through them contains one of the other $k-2$ points of $S^{\prime}$. The edge between $u$ and $v$ is drawn as the geodesic path $g(u,v)$. We claim that $G$ is planar, and the edges can be drawn without crossings inside $P$. Suppose there are two edges $g(u,v),g(x,y)$ in $G$, for $u,v,x,y\in S^{\prime}$, that intersect. By Lemma 2, one of the edges has the property that every geodesic disk through both of its endpoints leaves outside at least one of the endpoints of the other diagonal, which contradicts the definition of the edges of $G$. Thus, $G$ is a planar graph with $k$ vertices. By Euler’s formula, $|E|\leq 3k-6$, so $S^{\prime}$ contains at least $\binom{k}{2}-(3k-6)=\binom{k-3}{2}$ pairs that satisfy the property. ∎
@@
For any set $S$ of $n\geq 21$ points in a polygon, $\Pi^{in\mbox{-}out}(n)\geq\left\lceil\frac{16}{665}(n-2)\right\rceil\approx\left\lceil\frac{n-2}{41.6}\right\rceil\approx 0.024n$.
@@
In the same spirit as in the proof Theorem 2, we define two points in a subset $S^{\prime}\subset S$ of $k$ points to be dominated by $S^{\prime}$ if every geodesic disk through them leaves outside at least one of the other $k-2$ points of $S^{\prime}$.
Among any set of $k$ points in $P$ there are at most $\binom{k}{2}-\binom{k-3}{2}=3k-6$ pairs of points that do not satisfy the property of Lemma 6 and at most $63k-$ 6 pairs of points (possibly different than the previous pairs) that do not satisfy the property of Lemma 7. For $k\geq 11$, it holds that $2(3k-6)<\binom{k}{2}$, so that there is at least one pair of points $\{u,v\}\in S^{\prime}$ that satisfies both properties, hence dominates and is dominated by at least $\frac{\left[2\binom{k-3}{2}-\binom{k}{2}\right]\binom{n}{k}}{\binom{n}{2}}$ subsets of $S$. For each of these subsets, every geodesic disk through $u$ and $v$ has at least one of the other $k-2$ points of $S^{\prime}$, both inside and outside. Since a point of $S\setminus\{u,v\}$ could be in at most $\binom{n-3}{k-3}$ subsets containing $u$ and $v$, every geodesic disk through $u$ and $v$ has at least
$\frac{\left[2\binom{k-3}{2}-\binom{k}{2}\right]\binom{n}{k}}{\binom{n}{2}\binom{n-3}{k-3}}=(n-2)\frac{2(k-3)(k-4)-k(k-1)}{k(k-1)(k-2)}$
other points of $S$, both inside and outside. The largest coefficient of $n-2$ is obtained for $k=21$, thus $k=21$ leads to the largest possible coefficient, the one in the statement. ∎
@@
For any set $S$ of $n\geq 11$ points in a polygon, $\Pi^{in-out}(n)\geq\left\lceil\frac{n}{3+\sqrt{5}}\right\rceil+1\approx\left\lceil\frac{n}{5.2}\right\rceil\approx 0.1909n$.

### v2

@@ 4 The general case
In this section we provide a generalization to the geodesic setting of lower bounds that are known for the original problem in the plane. In particular, we generalize three results from [14, 13, 18]. We note that upper bounds for the Euclidean setting are in general also valid for the geodesic setting, since one can take the constructions provided in [2, 11, 14], and enclose the relevant set of points in a bounding polygon, big enough so that the geodesic disks through every three points coincide with the corresponding Euclidean disks.
4.1 A Constant Fraction of Points Inside Geodesic Disks
We begin by adapting some definitions and results from [18] to our setting. The intersection number $I(S)$ of a set of points $S$ in a polygon is defined as the number of different pairs of geodesic paths $g(u,v)$, $g(x,y)$ for $u,v,x,y\in S$ such that $g(u,v)$ intersects $g(x,y)$. Then, the intersection graph $G(S)=(V(G(S)),E(G(S)))$ is defined with $V(G(S))=\{g(u,v)\mid u,v\in S,u\neq v\}$, and two vertices $g(u,v)$, $g(x,y)$ being joined by an edge in $G(S)$ if $g(u,v)$ intersects $g(x,y)$ and $u,v,x,y$ are four distinct points. Finally, the oriented graph $\overrightarrow{G}(S)$ of $G(S)$ is defined orienting an edge $\{g(u,v),g(x,y)\}$ as $g(u,v)\rightarrow g(x,y)$ if any geodesic disk through $u$ and $v$ contains $x$ or $y$, or otherwise orienting $g(x,y)\rightarrow g(u,v)$. This orientation is consistent because of Lemma 2.
@@ 4.1 A Constant Fraction of Points Inside Geodesic Disks
For any $n\geq 5$, $\Pi(n)\geq\left\lceil\frac{n-2}{60}\right\rceil\approx 0.0166n$.
@@
By Kuratowski’s Theorem [16], each subset of $5$ points of a set $S$ contains four points $w,x,y,z$ such that $g(w,x)\cap g(y,z)\neq\emptyset$. Moreover, the subset $\{w,x,y,z\}$ appears in exactly n4n-4 subsets of $S$ with five elements. Thus, $I(S)\geq\frac{\binom{n}{5}}{n-4}$. By the definition of the intersection graph $G(S)$, $\lvert E(G(S))\rvert=I(S)$, thus $\lvert E(G(S))\rvert\geq\frac{\binom{n}{5}}{n-4}$.
@@
$\sum_{g(x,y)\in V(\overrightarrow{G}(S))}d^{+}(g(x,y))=\lvert E(G(S))\rvert\geq\frac{\binom{n}{5}}{n-4}.$
@@
This means that any geodesic disk through $u_{0}$ and $v_{0}$ contains at least one endpoint of $\frac{(n-2)(n-3)}{60}$ geodesic paths $g(x,y),\ x,y\in S$. Since each point could appear in at most n3n-3 of these geodesic paths, any geodesic disk through $u_{0}$ and $v_{0}$ contains at least $\left\lceil\frac{\frac{(n-2)(n-3)}{60}}{n-3}\right\rceil=\left\lceil\frac{n-2}{60}\right\rceil$ points of $S$ different from $u_{0}$ and $v_{0}$. ∎
@@
Among any set $S^{\prime}$ of $k$ points in a polygon $P$, there are at least $\binom{k-3}{2}$ pairs of points $x,y\in S^{\prime}$ such that every geodesic disk through $x$ and $y$ contains at least one of the other k2k-2 points of $S^{\prime}$.
@@
Let $G=(V,E)$ be the graph where $V\!=\!S^{\prime}$, and two vertices are joined by an edge if at least one of the geodesic disks through them does not contain any of the other k2k-2 points of $S^{\prime}$. The edge between $u$ and $v$ is drawn as the geodesic path $g(u,v)$. We claim that $G$ is planar, and the edges can be drawn without crossings inside $P$. Suppose there are two edges $g(u,v),g(x,y)$ in $G$, for $u,v,x,y\in S^{\prime}$, that intersect. By Lemma 2, one of the edges has the property that every geodesic disk through both of its endpoints contains at least one of the endpoints of the other edge, which contradicts the definition of the edges of $G$. Thus, $G$ is a planar graph with $k$ vertices. By Euler’s formula, $\lvert E\rvert\leq 3k-6$, so $S^{\prime}$ contains at least $\binom{k}{2}-(3k-6)=\binom{k-3}{2}$ pairs that satisfy the property. ∎
@@
For any $n\geq 8$, $\Pi(n)\geq\left\lceil\frac{5}{84}(n-2)\right\rceil=\left\lceil\frac{n-2}{16.8}\right\rceil\approx 0.0595n$.
@@
Two points in a subset $S^{\prime}\subset S$ of $k$ points dominate $S^{\prime}$ if every geodesic disk through them contains at least one of the other k2k-2 points of $S^{\prime}$. There are $\binom{n}{k}$ subsets of size $k$. By Lemma 6, each of these sets has $\binom{k-3}{2}$ dominating pairs. The average number of subsets dominated by a pair is $\frac{\binom{k-3}{2}\binom{n}{k}}{\binom{n}{2}}$. Thus, there is a pair of points $\{u,v\}\in S^{\prime}$ that dominates at least $\frac{\binom{k-3}{2}\binom{n}{k}}{\binom{n}{2}}$ subsets of $S$.
For each of these subsets, every geodesic disk through $u$ and $v$ contains at least one of the other k2k-2 points of $S^{\prime}$. Since a point of $S\setminus\{u,v\}$ can be in at most $\binom{n-3}{k-3}$ subsets that include $u$ and $v$, every geodesic disk through $u$ and $v$ contains at least
@@
other points of $S$. The largest coefficient of n2n-2 is $\frac{5}{84}$, obtained for $k=8$ and for $k=9$. ∎
We now present our strongest lower bound for $\Pi(n)$ by generalizing the approach by Edelsbrunner et al. [12] to the geodesic setting. This generalization is made by using order-$k$ geodesic Voronoi diagrams. An order-$k$ geodesic Voronoi diagram of a set $S$ in a polygon $P$ is the partitioning of $P$ into cells associated with all subsets $K\subseteq S$ of size $k$, such that each cell $V(K)$ is defined as $V(K)=\{x\in P\mid\max_{s\in K}g(x,s)\leq\min_{t\in S\setminus K}g(x,t)\}$, i.e., a partition of a polygon into regions such that all points in a region share the same $k$ geodesically nearest sites.
@@
For any $n\geq 6$, $\Pi(n)\geq\left\lceil\frac{n}{5}\right\rceil+1$.
@@
Unless specified otherwise, in this proof, order-$k$ diagrams refer to order-$k$ geodesic Voronoi diagrams. Recall that a geodesic disk through two points containing k1k-1 points in its interior has its center on an edge of the order-$k$ diagram. Consider the set of all disks that pass through two points $p,q\in S$ in the polygon. These disks are all centered on the bisector $b(p,q)$, and there are at most n2n-2 points on the bisector where a third point of $S$ is on the boundary of the disk. Since there is at most one geodesic disk through any three points, this partitions the bisector into at most n1n-1 pieces such that any disk through $p$ and $q$ centered on any point belonging to a fixed piece $s$ contains in its interior the same number of points $\rho(s)$. The pieces $s$ such that $\rho(s)=k-1$ are precisely the edges of the order-$k$ diagram.
@@
For some fixed bisector $b(p,q)$, let $b_{k}(p,q)$ be the union of the edges of the order-$1$ through order-$k$ diagrams on $b(p,q)$. The edges of $b_{k}(p,q)$ form a number of connected components, where a connected component is a maximal connected subset of a bisector that is the union of the closures of the edges of the order-$1$ through order-$k$ diagrams. Note that the number of connected components is strictly less than the total number of edges of all the diagram. Let $\lambda_{k}$ be the sum of the number of connected components formed by the edges of $b_{k}(p,q)$ over all pairs $p,q\in S$. The goal is to find the largest value of $k$ such that $\lambda_{k}<\binom{n}{2}$. Every bisector is completely covered by a single connected component when k=n2k=n-2. When $\lambda_{k}<\binom{n}{2}$, by the pigeonhole principle there exists a bisector $b(u,v)$ that does not have any edges of the order-$1$ through order-$k$ diagrams. Therefore, every geodesic disk with $u$ and $v$ on its boundary must contain in its interior at least $k$ points.
@@
We derive a formula for $\lambda_{k}$ in terms of the number of edges in the order-$1$ to the order-$k$ diagram by considering how the pieces must “pair up”, in the sense that a piece $s$ with $\rho(s)=i$ can only be adjacent to pieces $s^{\prime}$ of the same bisector with $\rho(s^{\prime})=i\pm 1$. If a piece $s$ with $\rho(s)=k-1$ is adjacent to two pieces $s_{1}$ and $s_{2}$ with $\rho(s_{1})=\rho(s_{2})=k-2$, the number of connected components is decreased by one. If $s$ is adjacent to only one piece $s_{1}$ with $\rho(s_{1})=k-2$, it extends an existing component. Otherwise, if it is not adjacent to any piece $s_{1}$ with $\rho(s_{1})=k-2$, it creates a new connected component. Let $\epsilon_{a},\epsilon_{b}$ and $\epsilon_{c}$ be, respectively, the number of pieces $s$ with $\rho(s)=k-1$ of the first, second, and third type. Then we have
$\lambda_{k}=\lambda_{k-1}+\epsilon_{c}-\epsilon_{a}.$ (1)
@@
Let $e_{k}$ be the number of edges in an order-$k$ diagram and let $S_{k}$ be the number of intersections between the order-$k$ diagram and the boundary of the polygon. The edges intersecting the boundary are analogous to the unbounded edges in the Euclidean setting. Therefore, in this proof we use the term endpoint to denote Voronoi vertices of the order-$k$ diagrams, not on the polygon boundary. Hence, the total number of endpoints of all order-$k$ edges is given by $2e_{k}-S_{k}$ since the edges that intersect the boundary of the polygon only contribute one endpoint. Moreover, we note that $e_{k}=\epsilon_{a}+\epsilon_{b}+\epsilon_{c}$.
@@
Let $\phi_{i}$ be the number of unpaired endpoints in $\lambda_{i}$. Consider all the unpaired endpoints of the connected components of $\lambda_{k-1}$. These endpoints can only come from edges in the order-k1{k-1} diagram since all edges of the order-$i$ diagrams for $i\in\{1,\ldots,k-2\}$ are all paired. Therefore, these unpaired endpoints can only pair up with edges from the order-$k$ diagram, which implies that $\phi_{k-1}=2\epsilon_{a}+\epsilon_{b}$. Thus, from Equation (1) and the fact that $e_{k}=\epsilon_{a}+\epsilon_{b}+\epsilon_{c}$, we have that $\lambda_{k}=\lambda_{k-1}+e_{k}-\phi_{k-1}$. If we iterate over $k$, we get that $\lambda_{k}=(e_{k}-\phi_{k-1})+(e_{k-1}-\phi_{k-2})+\ldots+(e_{2}-\phi_{1})+e_{1}=\sum_{i=1}^{k}e_{i}-\sum_{i=1}^{k-1}\phi_{i}$. Note that $\phi_{i-1}$ is the number of unpaired endpoints in $\lambda_{i-1}$, which are the paired endpoints of $\lambda_{i}$. Therefore, $\phi_{i}+\phi_{i-1}=2e_{i}-S_{i}$–the number of endpoints in order-$i$ diagrams. We use this to rewrite the equation for $\lambda_{k}$ in terms of $e_{i}$ and $S_{i}$.
[…]
@@
[…]
@@
We use the fact that $j(j+1)\leq\sum_{i=1}^{j}S_{i}\leq j(2n-j-1)$, see [8, Lemma $11$], which implies that
@@
which is a quadratic polynomial in $k$. We note that $\lambda_{k}<\binom{n}{2}$ whenever $k<\frac{n}{5}$. Thus if we set $k=\left\lceil\frac{n}{5}\right\rceil-1$, the pigeonhole principle implies that every geodesic disk with u{u} and v{v} on its boundary must contain in its interior at least $k$ points. Thus, we get that $\Pi(n)\geq{k+2}=\left\lceil\frac{n}{5}\right\rceil+1$, with the plus two accounting for u{u} and v{v}. ∎
@@ 4.2 A Constant Fraction of Points Inside and Outside Geodesic Disks
In this section we show that for a point set $S$, there is a pair of points $x,y\in S$ such that any disk through them contains a constant fraction of points of $S$, both inside and outside the disks. We obtain the first upper bound by modifying the previous proof of Theorem 2 by using Lemma 7. The result is analogous to Lemma 6 but, for completeness, we add here the proof of the bound.
@@
Among any set $S^{\prime}$ of $k$ points in a polygon, there are at least $\binom{k-3}{2}$ pairs of points $x,y\in S^{\prime}$ such that every geodesic disk through $x$ and $y$ leaves outside at least one of the other k2k-2 points of $S^{\prime}$.
@@
Let $G=(V,E)$ be the graph where $V\!=\!S^{\prime}$, and two vertices are joined by an edge if at least one of the geodesic disks through them contains all of the other k2k-2 points of $S^{\prime}$. The edge between $u$ and $v$ is drawn as the geodesic path $g(u,v)$. We claim that $G$ is planar, and the edges can be drawn without crossings inside $P$. Suppose there are two edges $g(u,v),g(x,y)$ in $G$, for $u,v,x,y\in S^{\prime}$, that intersect. By Lemma 3, one of the edges has the property that every geodesic disk through both of its endpoints leaves outside at least one of the endpoints of the other edge, which contradicts the definition of the edges of $G$. Thus, $G$ is a planar graph with $k$ vertices. By Euler’s formula, $\lvert E\rvert\leq 3k-6$, so $S^{\prime}$ contains at least $\binom{k}{2}-(3k-6)=\binom{k-3}{2}$ pairs that satisfy the property. ∎
@@
For any $n\geq 21$, $\Pi^{in\mbox{-}out}(n)\geq\left\lceil\frac{16}{665}(n-2)\right\rceil\approx\left\lceil\frac{n-2}{41.6}\right\rceil\approx 0.024n$.
@@
Let $S^{\prime}$ be a subset of $S$ of size $k$. Recall from the proof of Theorem 2 that we say two points of $S^{\prime}$ dominate $S^{\prime}$ if every geodesic disk through them contains at least one of the other k2k-2 points of $S^{\prime}$. In the same spirit, we define two points in $S^{\prime}$ to be dominated by $S^{\prime}$ if every geodesic disk through them leaves outside at least one of the other k2k-2 points of $S^{\prime}$.
There are $\binom{n}{k}$ subsets of size $k$ in $P$. Among any of these sets there are at most $\binom{k}{2}-\binom{k-3}{2}=3k-6$ pairs of points that do not satisfy the property of Lemma 6 and at most 3k63k-6 pairs of points (possibly different than the previous pairs) that do not satisfy the property of Lemma 7. Since there are $\binom{n}{2}$ pairs of points in $S$ and, for $k\geq 11$, it holds that $2(3k-6)<\binom{k}{2}$, there is at least one pair of points $u,v\in S^{\prime}$ that satisfies both properties, hence dominates and is dominated by at least $\frac{\left[\binom{k}{2}-2(3k-6)\right]\binom{n}{k}}{\binom{n}{2}}$} subsets of $S$. For each of these subsets, every geodesic disk through $u$ and $v$ has at least one of the other k2k-2 points of $S^{\prime}$, both inside and outside. Since a point of $S\setminus\{u,v\}$ could be in at most $\binom{n-3}{k-3}$ subsets containing $u$ and $v$, every geodesic disk through $u$ and $v$ has at least
$\frac{\left[\binom{k}{2}-2(3k-6)\right]\binom{n}{k}}{\binom{n}{2}\binom{n-3}{k-3}}=(n-2)\frac{2(k-3)(k-4)-k(k-1)}{k(k-1)(k-2)}$
other points of $S$, both inside and outside. The largest coefficient of $\frac{16}{665}$ for n2n-2 is obtained when $k=21$. ∎
Theorem 4 illustrates how the argument of [14] can be adapted to the geodesic setting. We improve the result by adapting the arguments from Theorem 3.
@@
For any $n\geq 14$, $\Pi^{in-out}(n)\geq\frac{7-\sqrt{37}}{12}n+O(1)\approx\frac{n}{13.08}$.